JIz Instructions Question 6 10 pts Suppose Sues monthly demand equation for coffee is Q = 30 - 3P, and the price is $4/cup.
Answer #1

The Alt BR JOA le Ps4 Where AB = o 15 18 30 (a) Q=30-3P At P= $4, Q = 30 - 3 (4) > Q = 18 So, Sue will buy 18 cups in one mon